how to solve x-y=11, 2x+y=19 in substitution method
Answer by checkley79(3341)

x-y=11 or x=y+11
2x+y=19
2(y+11)+y=19
2y+22+y=19
3y=19-22
3y=-3
y=-1 ans.
x=-1+11
x=10 ans.
Proof:
2*10-1=19
20-1=19
19=19